# Circumgalactic Enrichment by a Multi-Galaxy Merger at Redshift 6.7

### v1

@@ Methods
JWST Spectroscopy data
@@
The spectroscopic data of ELG1 and ELG5 used in this work were also obtained as part of the JADES NIRSpec observations. The JADES program opened three adjacent shutters for each target and nodded between these shutters (perpendicular to the dispersion direction) to improve the background subtraction and bad pixel rejection. JADES obtained low-resolution spectra () using the NIRSpec disperser/filter configuration of PRISM/CLEAR, which provides wavelength coverage from 0.6 to 5.3 m. The exposure times of ELG1 and ELG5 were 2.1 and 9.3 hours, respectively. JADES also obtained medium-resolution spectra () using the disperser/filter configurations of G140M/F100LP, G140M/F070LP, and G235M/F170LP, G395M/F290LP, which provides a full wavelength coverage from 0.7 to 5.3 m. The exposure times of ELG1 for the three configurations were 3.4, 2.8, and 2.8 hours, and the exposure times of ELG5 for the three configurations were 2.3 hours. The JADES NIRSpec data used in this work was downloaded from the MAST HLSP archive111https://archive.stsci.edu/hlsp/jades. Please refer to ref [58] for details.
Photometric redshifts of ELGs
@@
Based on the JADES photometry, we use the Python package Bagpipes [59] to perform SED fitting including measuring redshifts of ELG2, ELG3, and ELG4. We employ Binary Population and Spectral Synthesis v2.2.1 [60, 61] (BPASS) with a broken power-law initial mass function with slopes of for stars with 0.1 – 0.5 and for 0.5 – 100 (model ‘135_100’). For the star formation history parameterization, we use the Gaussian Process model from DENSEBASIS [62], where the star formation history is split into 4 dynamically adjusted time bins, and during each time bin, of the total stellar mass is formed. The metallicity is allowed to span 0.001 – 1 . We adopt the Calzetti [63] dust attenuation law with ranging from 0.0 to 4.0 mag. We include nebular emission with the metallicity of the gas equal to that of the stellar populations, and an ionization parameter, in the range to . We use the uniform priors for the Bagpipes parameters across the allowed range. For the redshift, we also adopt a uniform prior across the range of 5 – 8.5. Figure 2 shows the probability distribution function of the redshifts and the best-fit models for ELG2, ELG3, and ELG4.
@@
We also use Bagpipes to estimate the slitlosses of NIRSpec spectroscopy of ELG1 and ELG5. We adopt a multiplicative scaling factor (assumed to be a second-order Chebyshev polynomial) in Bagpipes to correct the wavelength-dependent spectral flux calibration to the broadband photometry. We allow the zeroth order to vary from 0.1 – 10 because of the relatively large slitloss in ELG1, while the first and second order are allowed to vary from -0.5 – 0.5. In Figure 2, we present the flux-calibrated spectra of ELG1 and ELG5.
| Galaxy | clump | SFR10Myr | MUV | ||||
| ID | ID | yr-1 | [erg s-1] | pkpc | |||
| ELG1 | C1 | ||||||
| C2 | |||||||
| C3 | |||||||
| C4 | |||||||
| ELG2 | C5 | ||||||
| C6 | |||||||
| C7 | |||||||
| ELG3 | C8-1 | ||||||
| C8-2 | |||||||
| C9 | |||||||
| ELG4 | C10 | ||||||
| ELG5 | C11 | ||||||
| C12 | |||||||
| C13 | |||||||
| C14 | |||||||
| C15 | |||||||
| C16 | – | ||||||
@@
Clump identification and photometry
Our clump identification method is illustrated in the Extended Data Figure 1. We identify the clumps from the NIRCam F115W image, as it offers the highest spatial resolution compared to the NIRCam images at longer wavelengths and is unaffected by the Ly break, unlike F090W. We first extract a cutout, corresponding to a physical size of pkpc2. To suppress the outshining components and enhance the visibility of clump cores, we smooth the cutout using a two-dimensional Gaussian kernel ( 1 pixel; 0.03 arcsec) and subtract the smoothed image from the original F115W cutout to produce a contrast image. We then use the Python Package sep [64, 65] to detect clumps in the contrast image and extract their positions. The red circles in Extended Data Figure 1c,d mark the identified clumps. Additionally, we notice several clumps that are clearly detected in the F356W image, but are only marginally detected in F115W. These clumps are manually added to the catalog based on their positions in the F356W image, and they are marked as green circles in Extended Data Figure 1d.
@@
We first perform the morphological fitting to the F115W image. We adopt the uniform priors for the Sérsic parameters and the Gaussian priors for the positions with the mean to be those measured by sep. Extended Data Figure 1 panels and show the best-fit model and the residual map of F115W, respectively. After obtaining the best-fit parameters from F115W, we use these measurements as the priors for fitting the clumps in other bands. Due to the substantial PSF variation across JWST images (from in F115W to in F444W), deblending adjacent clumps becomes increasingly challenging at longer wavelengths. Therefore, to address this, we fix the pixel centers and Sérsic parameters to the values obtained from F115W, and only the fluxes are allowed to vary when fitting the other bands (F090W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W).
Photometric redshifts and physical properties of clumps
@@
Using the photometry described in the previous section, we perform model fitting for the clumps again using Bagpipes. This step is necessary to identify the clumps that are associated with JQ and remove other foreground clumps unassociated with the JQ galaxies. The Bagpipes configuration is identical to the one used for the photometric redshift estimation of ELG2, ELG3, and ELG4. We visually inspect the best-fit results and define a clump as a member of JQ if its best-fit photometric redshift lies within the range of 6.65 and 6.75. Once the member clumps of JQ are identified, we rerun the SED fitting with the redshift fixed to 6.71 to best model their properties assuming the spectroscopic redshift of ELG1 and ELG5. We note clumps C8-1 and C8-2 are heavily blended in the longer wavelength imaging data, so we combine their photometry for the SED fitting process. In Extended Data Table 1, we summarize the physical properties of the identified member clumps.
@@
In Extended Data Figure 2, we compare the size – SFR relation of the clumps in the JQ system with those of the galaxies at similar redshifts observed with JWST NIRCam [22, 21, 23, 24]. Several studies [22, 21, 23] resolved the galaxies through the lensing technique and found that the star clusters and star-forming clumps identified within the individual galaxies have sizes pc and SFRs yr-1. Their sizes and SFRs are significantly smaller than those of the clumps in the JQ system. The JQ clumps have sizes and SFRs more similar to those unresolved galaxies at 5 – 14 [24], which typically have sizes pc and SFRs yr-1.
[O III]+H emission map of JQ
@@
Spatially-resolved photometry and properties
@@
As aforementioned, we fix the half-light radii of clumps when fitting their morphologies in the F356W image. However, emission lines are frequently observed to be more spatially extended than the underlying stellar continuum [68], so this approach could miss diffuse [O III]+H emission from the clumps, underestimation of their flux densities. Moreover, the simple Sérsic profile is not always able to accurately describe the morphologies of the clumps due to their disturbed and irregular shapes. To address these limitations and provide a robust estimate of the total stellar mass and SFR of the JQ system, we perform spatially resolved SED fitting directly using multi-wavelength images. This method is insensitive to the assumption of morphologies across different wavelengths and has been extensively validated in previous studies [69]. For this analysis, we adopt the SED fitting package, CIGALE [70], which offers computational efficiency when fitting a large number of data points. This also allows us to assess the potential biases associated with the SED fitting methods.
@@
Our method follows the procedure in ref [69]. In detail, we used the PSF-matched imaging with a pixel scale of 0.03 arcsec. To enhance the SNR of regions surrounding and connecting the galaxies in the JQ, we applied an adaptive binning algorithm222https://github.com/pierrethx/MVT-binning based on the Weighted Voronoi Tessellation method described in ref [71, 72]. We chose a minimum SNR of 8 in the F356W band, and a minimum Voronoi bin size of 9 pixels, which is approximately the size of the FWHM of the PSF of F356W (FWHM = 01).
@@
We run CIGALE using the 9-band fluxes in each Voronoi bin. Here we described the modules and parameters adopted in CIGALE. We adopted a delayed SFH allowing the and stellar age to vary from 0.05 – 1 Gyr and 40 – 800 Myr, respectively. In addition to the main SFH, we allow for a more recent “burst” of star formation as an exponential SFH with a allowed to vary from 20 – 100 Myr, an age of 1 – 30 Myr, and a mass fraction of this burst () to vary from 0 – 0.9. We assumed a ref [73] IMF and the stellar population synthesis models presented by ref [74] with metallicity ranging from Solar () to sub-Solar values (0.4 and 0.2 ). We include nebular emission using templates of ref [75]. We allow the ionization parameter to vary between to , the gas metallicity () to vary between 0.002 and 0.02, and a fixed electron density of 100 cm-3. The dust attenuation follows the Calzetti extinction law [63] allowing the dust attenuation in emission lines from nebular regions to vary from 0 to 1.1, and a lower dust attenuation in stellar continuum with a fixed dust attenuation ratio () between emission lines and stellar continuum. In Extended Data Figure 4, we present the spatially resolved maps of mass and SFR.
@@
The integrated properties of the JQ system are derived by summing the spatially resolved maps. We obtain a total stellar mass of and a total SFR averaged of yr-1. We estimate the uncertainties in the total stellar mass and SFR by propagating the measurement errors from each Voronoi bin. The total stellar mass is consistent with the sum of the stellar masses of individual clumps from Bagpipes, however, the total SFR is nearly double the sum of the SFRs of the clumps (see Extended Data Table 1). To investigate this discrepancy, we further use CIGALE to model the SEDs of 17 clumps and find that the estimated stellar masses and SFRs agree with those derived from Bagpipes within uncertainties. Therefore, the difference between the SFRs estimated from the spatially resolved map and the individual clumps is not due to the choice of the SED fitting method. These results imply that while the majority of stellar mass is formed within the clumps, a large fraction, , of the total SFR may arise from diffuse emission surrounding the clumps if attributing the diffuse emission to star formation.
Contextualizing JQ-like systems with cosmological simulations
@@
We compare the presence and properties of the JQ system with galaxy-size clumps found in a simulated lightcone [26]. The simulated lightcone spans 2 deg2, containing simulated galaxies arranged in a way that mimics how we observe the real Universe. The dataset provides galaxy coordinates (RA and Dec) on the projected plane, and includes both cosmological redshifts and peculiar velocities of galaxies along the line of sight. This enables a direct comparison between the simulation and observation. The large volume of the lightcone enables more precise quantification of the number of JQ-like systems in the Universe, although this is significantly larger than the 120 arcmin2 covered by the JADES dataset. The lightcone is constructed by drawing halos from the SMDPL simulation from the MultiDark suite [76] at various redshifts based on their line-of-sight distances. Within these dark matter halos, galaxies are then simulated using the Santa Cruz semi-analytic model for galaxy formation [77, 78, 79, 80], which incorporates the evolution of a wide variety of physical processes, such as cosmological accretion, cooling, star formation, chemical enrichment, and stellar and AGN feedback. The performance of the model at high redshift is validated against one-point distribution functions of , , and SFR [79, 81] and galaxy clustering constraints [80]. We also note that dust attenuation is accounted for in the predicted photometry of simulated galaxies. Sky coordinates of the halos and galaxies therein are computed based on the physical coordinates of sources within the lightcone. We refer the readers to refs. [80, 26] for detailed descriptions for the simulation and the construction of the lightcone.
@@
While the Santa Cruz SAM tracks the decay of satellite orbits due to dynamical friction and the tidal disruption of infalling satellites (see ref [82], for details), the model does not explicitly model star-forming clumps nor clumps that form during mergers. For comparison with the lightcone, we only consider the galaxy-size clumps in JQ that are distinctly separated to be the analogs to the galaxies in the lightcone. These include ELG1 (C1+C2+C3), ELG2 (C5+C6), ELG3 (C8+C9), ELG4 (C10), ELG5 (C11), C4, C7, C12+C13, C14, C16. We exclude C15 because it is relatively faint compared to the others. From the lightcone, we first create a parent galaxy sample at – 7.0 with F277W magnitude and then identify a subset of galaxy systems in the parent sample by requiring at least one companion with a velocity difference of km s-1 and a projected separation of arcsec. These criteria are chosen to be consistent with the observation of JQ and the depth of the JADES survey. The redshift range (6.5 – 7.0) corresponds to a comoving volume of cMpc3 for the lightcone simulation. In total, we select 29,696 unique systems (central and satellite galaxies) from the 2 deg2 lightcone, of which 3154 systems have at least two galaxy-size clumps (inferred by simulated effective radii and SFRs based on results presented in Ref [24]). This implies that we would expect galaxy systems with galaxy-size clumps at – 7.6 for the JADES survey area. In Extended Data Figure 5, we present the expected distribution of clump numbers in galaxy systems in the JADES survey. Using this criterion (systems with clumps), we identify 8 such systems in the 2 deg2 lightcone, corresponding to an expectation of just 0.26 systems within the JADES survey volume. This indicates the galaxy-size clump density in JQ is higher than galaxy systems in cosmological simulation.
@@
The JQ system is serendipitously discovered in an ELG study based on the JADES F410M image. Given the relatively small survey volume of the JADES F410M image ( arcsec2), we estimate the observed comoving number density of the JQ system to be cMpc-3. The error is estimated based on the Poisson error of the small number statistic [83]. In comparison, we estimate the comoving number density of JQ-like systems in lightcone simulation to be cMpc-3.
Merging timescale of the JQ system

### v2

@@ 1.1 HST and JWST Imaging data
1.2 JWST Spectroscopy data
@@
The spectroscopic data of ELG1 and ELG5 used in this work were also obtained as part of the JADES NIRSpec observations. The JADES program opened three adjacent shutters for each target and nodded between these shutters (perpendicular to the dispersion direction) to improve the background subtraction and bad pixel rejection. JADES obtained low-resolution spectra () using the NIRSpec disperser/filter configuration of PRISM/CLEAR, which provides wavelength coverage from 0.6 to 5.3 m. The exposure times of ELG1 and ELG5 were 2.1 and 9.3 hours, respectively. JADES also obtained medium-resolution spectra () using the disperser/filter configurations of G140M/F100LP, G140M/F070LP, and G235M/F170LP, G395M/F290LP, which provides a full wavelength coverage from 0.7 to 5.3 m. The exposure times of ELG1 for the three configurations were 3.4, 2.8, and 2.8 hours, and the exposure times of ELG5 for the three configurations were 2.3 hours. The JADES NIRSpec data used in this work were downloaded from the MAST HLSP archive111https://archive.stsci.edu/hlsp/jades. Please refer to ref [61] for details.
1.3 Photometric redshifts of ELGs
@@
Based on the JADES photometry, we use the Python package Bagpipes [62] to perform SED fitting, including measuring redshifts of ELG2, ELG3, and ELG4. We note that although the JWST images show ELG3 and ELG4 likely connected, we consider them as two separate galaxies because of their distinct stellar populations. As shown in Figure 3 and Extended Data Table 1, ELG3 is the most massive galaxy with a moderate SFR consistent with being located on the SFMS. In contrast, ELG4 has significantly elevated SFR, placing it above SFMS by dex. This indicates ELG3 is dominated by more mature populations, but ELG4 is dominated by newly born stars. We employ Binary Population and Spectral Synthesis v2.2.1 [63, 64] (BPASS) with a broken power-law initial mass function with slopes of for stars with 0.1 – 0.5 and for 0.5 – 100 (model ‘135_100’). For the star formation history parameterization, we use the Gaussian Process model from dense_basis [65], where the star formation history is split into 4 dynamically adjusted time bins, and during each time bin, of the total stellar mass is formed. The metallicity is allowed to span 0.001 – 1 . We adopt the Calzetti [66] dust attenuation law with ranging from 0.0 to 4.0 mag. We include nebular emission with the metallicity of the gas equal to that of the stellar populations, and an ionization parameter, in the range to . We use the uniform priors for the Bagpipes parameters across the allowed range. For the redshift, we also adopt a uniform prior across the range of 0 – 8.5. Figure 2 shows the probability distribution function of the redshifts and the best-fit models for ELG2, ELG3, and ELG4.
@@
We also use Bagpipes to estimate the slitlosses of NIRSpec spectroscopy of ELG1 and ELG5. We adopt a multiplicative scaling factor (assumed to be a second-order Chebyshev polynomial) in Bagpipes to correct the wavelength-dependent spectral flux calibration to the broadband photometry. We allow the zeroth order to vary from 0.1 – 10 because of the relatively large slitloss in ELG1, while the first and second orders are allowed to vary from -0.5 – 0.5. In Figure 2, we present the flux-calibrated spectra of ELG1 and ELG5.
| Galaxy | clump | RA | DEC | SFR10Myr | MUV | ||||
| ID | ID | yr-1 | [erg s-1] | pkpc | |||||
| ELG1 | C1 | 3:32:37.19 | -27:48:57.04 | ||||||
| C2 | 3:32:37.19 | -27:48:56.83 | |||||||
| C3 | 3:32:37.18 | -27:48:56.76 | |||||||
| C4 | 3:32:37.19 | -27:48:55.75 | |||||||
| ELG2 | C5 | 3:32:37.21 | -27:48:55.50 | ||||||
| C6 | 3:32:37.21 | -27:48:55.26 | |||||||
| C7 | 3:32:37.19 | -27:48:55.03 | |||||||
| ELG3 | C8-1 | 3:32:37.30 | -27:48:54.92 | ||||||
| C8-2 | 3:32:37.30 | -27:48:55.02 | |||||||
| C9 | 3:32:37.31 | -27:48:55.03 | |||||||
| ELG4 | C10 | 3:32:37.28 | -27:48:54.83 | ||||||
| ELG5 | C11 | 3:32:37.39 | -27:48:54.75 | ||||||
| C12 | 3:32:37.19 | -27:48:56.20 | |||||||
| C13 | 3:32:37.21 | -27:48:56.33 | |||||||
| C14 | 3:32:37.17 | -27:48:55.28 | |||||||
| C15 | 3:32:37.41 | -27:48:54.24 | |||||||
| C16 | 3:32:37.13 | -27:48:56.19 | – | ||||||
@@
1.4 Clump identification and photometry
Our clump identification method is illustrated in the Extended Data Figure 1. We identify the clumps from the NIRCam F115W image, as it offers the highest spatial resolution compared to the NIRCam images at longer wavelengths and is unaffected by the Ly break, unlike F090W. We first extract a cutout, corresponding to a physical size of pkpc2. To suppress the outshining components and enhance the visibility of clump cores, we smooth the cutout using a two-dimensional Gaussian kernel and subtract the smoothed image from the original F115W cutout to produce a contrast image. A kernel that is too narrow (e.g., pixel) excessively subtracts flux from the clump cores, making them difficult to detect, while a kernel that is too broad (e.g., pixel) can produce a large dark halo surrounding the central bright clump, resulting in contamination to its nearby objects. To balance between these two effects, we adopt a kernel size of pixel (0.06 arcsec), of which the FWHM is comparable to the separation between close clumps in JQ. This configuration effectively preserves the clump cores while minimizing contamination to neighboring sources. We then use the Python Package sep [67, 68] to detect clumps in the contrast image and extract their positions. The red circles in Extended Data Figure 1c,d mark the automatically identified clumps. However, as F115W traces the rest-frame UV emission of objects, which is sensitive to dust attenuation, this method could miss clumps that are heavily obscured. In contrast, the F356W image probes the rest-frame optical emission of objects, which is less affected by dust and can be significantly enhanced by strong [O III]+H emission. Therefore, F356W provides a valuable complement for identifying clumps that are attenuated. Through a visual inspection of the F356W image, we identify 5 additional clumps that are missed in the F115W contrast image. These clumps are then manually added to the catalog based on their positions in the F356W image, and they are marked as green circles in Extended Data Figure 1d. In total, we identify 17 clumps in JQ: 3 in ELG1, 2 in ELG2, 3 in ELG3, 1 each in ELG4 and ELG5, and 7 additional clumps in the surrounding region. We acknowledge that the spatially inhomogeneous dust distribution could artificially divide a clump into multiple clumps, leading to an overestimation of the number of clumps. Confirming this would require a high-resolution, dust-insensitive observation, such as ALMA [C II] observations. Nonetheless, this does not affect our subsequent comparison with simulations, because we consider the close clumps as a single system in the analysis.
@@ 1.4 Clump identification and photometry
We first perform the morphological fitting to the F115W image. We adopt the uniform priors for the Sérsic parameters and the Gaussian priors for the positions with the mean to be those measured by sep. For the fluxes, we adopted the Gaussian priors with the mean of 0 Jy and the standard deviation of 0.2 Jy to cover all the possible ranges. Extended Data Figure 1 panels and show the best-fit model and the residual map of F115W, respectively. The residual map of F115W does not show clear large structure, indicating that we have identified all the bright components and well-modeled their lights. After obtaining the best-fit parameters from F115W, we use these measurements as the priors for fitting the clumps in other bands. Due to the substantial PSF variation across JWST images (from in F115W to in F444W), deblending adjacent clumps becomes increasingly challenging at longer wavelengths. Therefore, to address this, we fix the pixel centers and Sérsic parameters to the values obtained from F115W, and only the fluxes are allowed to vary when fitting the other bands (F090W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W).
1.5 Photometric redshifts and physical properties of clumps
@@
Using the photometry described in the previous section, we perform model fitting for the clumps again using Bagpipes. This step is necessary to identify the clumps that are associated with JQ and remove other foreground clumps unassociated with the JQ galaxies. The Bagpipes configuration is identical to the one used for the photometric redshift estimation of ELG2, ELG3, and ELG4. We visually inspect the best-fit results and define a clump as a member of JQ if its best-fit photometric redshift lies within the range of 6.65 and 6.75. Once the member clumps of JQ are identified, we rerun the SED fitting with the redshift fixed to 6.71 to best model their properties assuming the spectroscopic redshift of ELG1 and ELG5. We note that clumps C8-1 and C8-2 are heavily blended in the longer wavelength imaging data, so we combine their photometry for the SED fitting process. In Extended Data Table 1, we summarize the physical properties of the identified member clumps.
In Figure 3 and Extended Data Figure 2, we compare the SFRs of the JQ galaxies and clumps with the galaxies and the SFMS at similar redshift from the literature [27, 29, 30, 28]. Most of the galaxies and clumps in JQ show elevated SFRs compared to the median SFMS relation. Recently, it has been recognized that the high-redshift galaxies experience bursty star formation [71], where a low-mass galaxy in the bursty phase can significantly increase its opportunity of being detected. However, this observational bias is unlikely to affect our analysis of the JQ clumps. JQ is in the deepest region of the JADES survey, which provides sufficient sensitivity to detect a complete low-mass sample reaching . Specifically, the purple diamonds in Extended Data Figure 2 are the spectroscopically observed galaxies selected from the JADES survey [29]. This sample includes a large number of galaxies with stellar masses comparable to those of the JQ clumps, but with much lower SFRs. Therefore, the elevated SFRs observed in the JQ clumps are intrinsic rather than due to this selection effects.
@@
In Extended Data Figure 3, we compare the size – SFR relation and size – mass relation of the clumps in the JQ system with those of the galaxies at similar redshifts observed with JWST NIRCam [21, 20, 22, 23, 72, 73]. Several studies [21, 20, 22, 72] resolved the galaxies through the lensing technique and found that the star clusters and star-forming clumps identified within the individual galaxies have sizes pc and SFRs yr-1. Their sizes and SFRs are significantly smaller than those of the clumps in the JQ system. The JQ clumps have sizes and SFRs more similar to those unresolved galaxies at 5 – 14 [23], which typically have sizes pc and SFRs yr-1. In addition, as shown in Extended Data Figure 3, the JQ clumps generally follow the size – mass relation measured for galaxies, whereas the lensing-resolved clumps in previous studies have much smaller sizes at a fixed mass. Therefore, to distinguish them from these smaller structures, we refer to the JQ clumps as galaxy-size clumps. Given the limited resolution of JWST observations, we do not attempt to further classify the clumps into morphological components such as bulges and disks in this work.
1.6 [O III]+H emission map of JQ
@@
1.7 Spatially-resolved photometry and properties
@@
As aforementioned, we fix the half-light radii of clumps when fitting their morphologies in the F356W image. However, emission lines are frequently observed to be more spatially extended than the underlying stellar continuum [74], so this approach could miss diffuse [O III]+H emission from the clumps, leading to an underestimation of their flux densities. The simple Sérsic profile is not always able to accurately describe the morphologies of the clumps due to their disturbed and irregular shapes. Moreover, when galaxies show high specific star formation, the young, luminous stellar populations can outshine the older populations and dominate the light, resulting in further underestimation of stellar mass [75]. To address these limitations and provide a robust estimate of the total stellar mass and SFR of the JQ system, we perform spatially resolved SED fitting directly using multi-wavelength images. This method is insensitive to the assumption of morphologies across different wavelengths and has been extensively validated in previous studies [76]. For this analysis, we adopt the SED fitting package, CIGALE [77], which offers computational efficiency when fitting a large number of data points. This also allows us to assess the potential biases associated with the SED fitting methods.
@@
Our method follows the procedure in ref [76]. In detail, we used the PSF-matched images (to F444W) with a pixel scale of 0.03 arcsec. To enhance the SNR of regions surrounding and connecting the galaxies in the JQ, we applied an adaptive binning algorithm based on the Weighted Voronoi Tessellation method described in ref [78, 79]. We chose a minimum SNR of 8 in the F356W band, and a minimum Voronoi bin size of 9 pixels, which is approximately the size of the FWHM of the PSF of F356W (FWHM = 01). It has been recognized that, with increasing spatial resolution, the smallest resolved elements may have too little stellar mass for the stellar initial mass function (IMF) to be fully populated. This effect, known as stochastic IMF sampling, can bias the derived properties such as stellar mass, age, and SFR. Our binning procedure ensures that the typical stellar masses of the resulting bins are , which is well above the threshold (; e.g., [80]) where stochastic IMF sampling is expected to significantly affect results. Therefore, this effect does not impact our analysis.
@@
We run CIGALE using the 9-band fluxes in each Voronoi bin. Here we describe the modules and parameters adopted in CIGALE. We adopt a delayed SFH allowing the and stellar age to vary from 0.05 – 1 Gyr and 40 – 800 Myr, respectively. In addition to the main SFH, we allow for a more recent “burst” of star formation as an exponential SFH with a allowed to vary from 20 – 100 Myr, an age of 1 – 30 Myr, and a mass fraction of this burst () to vary from 0 – 0.9. We assume a ref [81] IMF and the stellar population synthesis models presented by ref [82] with metallicity ranging from Solar () to sub-Solar values (0.4 and 0.2 ). We include nebular emission using templates of ref [83]. We allow the ionization parameter to vary between to , the gas metallicity () to vary between 0.002 and 0.02, and a fixed electron density of 100 cm-3. The dust attenuation follows the Calzetti extinction law [66] allowing the dust attenuation in emission lines from nebular regions to vary from 0 to 1.1, and a lower dust attenuation in stellar continuum with a fixed dust attenuation ratio () between emission lines and stellar continuum. In Extended Data Figure 5, we present the spatially resolved maps of mass and SFR.
@@
The integrated properties of the JQ system are derived by summing the spatially resolved maps. We obtain a total stellar mass of and a total SFR averaged of yr-1. We estimate the uncertainties in the total stellar mass and SFR by propagating the measurement errors from each Voronoi bin. The total stellar mass is consistent with the sum of the stellar masses of individual clumps from Bagpipes. This consistency implies that we do not identify the stellar population variation at a clump-size scale, and the majority of stellar mass is formed within the clumps. However, the total SFR is nearly double the sum of the SFRs of the clumps (see Extended Data Table 1). To investigate this discrepancy, we use CIGALE to model the SEDs of 17 clumps and find that the estimated stellar masses and SFRs agree with those derived from Bagpipes within uncertainties. Therefore, the difference between the SFRs estimated from the spatially resolved map and the individual clumps is not due to the choice of the SED fitting method. Instead, the difference in the derived SFRs suggests that a large fraction, , of the total SFR may arise from diffuse emission surrounding the clumps if attributing the diffuse emission to star formation.
1.8 Contextualizing JQ-like systems with cosmological simulations
@@
We compare the presence and properties of the JQ system with galaxy-size clumps found in a simulated lightcone [26]. The simulated lightcone spans 2 deg2, containing simulated galaxies arranged in a way that mimics how we observe the real Universe. The dataset provides galaxy coordinates (RA and Dec) on the projected plane, and includes both cosmological redshifts and peculiar velocities of galaxies along the line of sight. This enables a direct comparison between the simulation and observation. The large volume of the lightcone enables more precise quantification of the number of JQ-like systems in the Universe, although this is significantly larger than the 120 arcmin2 covered by the JADES dataset. The lightcone is constructed by drawing halos from the SMDPL simulation from the MultiDark suite [84] at various redshifts based on their line-of-sight distances. Within these dark matter halos, galaxies are then simulated using the Santa Cruz semi-analytic model (SC SAM) for galaxy formation [85, 86, 87, 88], which incorporates the evolution of a wide variety of physical processes, such as cosmological accretion, cooling, star formation, chemical enrichment, and stellar and AGN feedback. The SAM approach does not rely on spatially resolved elements like mass particles and grid cells, but rather utilises analytical and empirical recipes to model the formation and growth of galaxies within halos and their merger histories. Dark matter halos in the lightcone are well-resolved and complete down to , which approximately corresponds to a stellar mass of . The performance of the model at high redshift is validated against one-point distribution functions of , , and SFR [87, 89] and galaxy clustering constraints [88]. We also note that dust attenuation is accounted for in the predicted photometry of simulated galaxies. However, the lightcone simulation does not employ a prescription for emission lines, which limits us to directly comparing the emission line properties of simulated galaxies and JQ galaxies. Sky coordinates of the halos and galaxies therein are computed based on the physical coordinates of sources within the lightcone. We refer the readers to refs. [88, 26] for detailed descriptions for the simulation and the construction of the lightcone.
@@
From the lightcone, we first create a parent galaxy sample at – 7.0 with F277W magnitude and then identify a subset of galaxy systems in the parent sample by requiring at least one companion with a velocity difference of km s-1 and a projected separation of arcsec. These criteria are chosen to be consistent with the observation of JQ and the depth of the JADES survey. The redshift range (6.5 – 7.0) corresponds to a comoving volume of cMpc3 for the lightcone simulation. In total, we select 29,696 unique systems (central and satellite galaxies) from the 2 deg2 lightcone, of which 3154 systems have at least two galaxies. This implies that we would expect galaxy systems with galaxies at – 7.6 for the JADES survey area. In Extended Data Figure 6, we present the expected distribution of clump numbers in galaxy systems in the JADES survey.
The SAM tracks the decay of satellite orbits due to dynamical friction and the tidal disruption of infalling satellites (see ref [90], for details), which is capable of reproducing separations that are comparable to the JQ galaxies and clumps. However, as the SAM approach treats different species of baryonic matter (e.g., hot gas, cold gas, stars) as reservoirs and relies on analytic recipes to track the conversion of mass between different reservoirs, the SAM does not spatially resolve individual star-forming clumps within a star-forming disk or model the clumps that are formed through the merger-induced star formation (for example, merger debris). On the other hand, due to the lack of kinematic information of JQ clumps, we cannot distinguish the physical origins of these clumps: some clumps may result from merger-induced star formation, which would not be tracked by the SAM and therefore should be excluded in comparison; others may have formed within companion dark matter halos before merger, in which case they should be included. Thus, a careful assessment of the number of galaxies in JQ is critical for a fair comparison with the lightcone.
We first consider the galaxy systems and their clumps in JQ that are separated to be the analogs to the galaxies in the lightcone. Clumps that are closely spaced are considered star-forming regions within the same galaxy. These include ELG1 (C1+C2+C3), ELG2 (C5+C6), ELG3 (C8+C9), ELG4 (C10), ELG5 (C11), C4, C7, C12+C13, C14, C16. We exclude C15 because it is relatively faint compared to the others, and again, we treat ELG3 and ELG4 as separate galaxies due to their distinct stellar populations. Using this definition (systems with clumps), we identify 8 such systems in the 2 deg2 lightcone, indicating that JQ has a number density larger than of galaxy systems in the simulation. Among these 8 simulated systems, we further select major mergers by requiring a stellar mass ratio between the primary and secondary galaxies to be less than 4 [91, 92]. This yields only 4 major merger systems in the lightcone, corresponding to an expected number of just 0.13 such systems within the JADES survey volume.
To consider a more conservative case, we restrict our comparison to six clearly isolated galaxies/clumps: ELG1 (C1+C2+C3), ELG2 (C5+C6), ELG3 (C8+C9), ELG4 (C10), ELG5 (C11), and C16. Here, we exclude the clumps surrounding the five main galaxies to test whether the lack of JQ-like systems in the simulation may be due to the omission of clumps formed during merger events. In this case, we identify 28 systems with galaxies in the 2 deg2 lightcone, and 11 of which are major merger systems. This increases the expected number of JQ-like systems within the JADES survey volume to 0.36, but is still below the observed density.
These comparisons indicate that the JQ-like systems appear to be under-represented by the simulated lightcone. In the SC SAM framework, once a galaxy becomes a satellite, its hot gas supply from cosmological accretion is immediately shut off (under the assumption that the hot gas is added to the reservoir of its respective central galaxy), effectively cutting off the supply of new fuel and causing it to cease star formation prematurely. Together with the lack of explicitly modeled star-forming clumps, these scenarios may be responsible for this underestimation. A recent study comparing the observed galaxy pair fraction with predictions from the SC SAM simulation also found that the simulation underpredicts the pair fraction at [91], which is consistent with this interpretation.
@@
The JQ system was serendipitously discovered in an ELG study based on the JADES F410M image. Given the relatively small survey volume of the JADES F410M image ( arcsec2), we estimate the observed comoving number density of the JQ system to be cMpc-3. The error is estimated based on the Poisson error of the small number statistic [93]. In comparison, we estimate the comoving number density of JQ-like systems in lightcone simulation to be cMpc-3.
1.9 Merging timescale of the JQ system